OpenStudy (anonymous):

02:35:55 Convert each rate using dimensional analysis. Round to the nearest tenth, if necessary. 30 mi / h = ____ ft / min

OpenStudy (anonymous):

How many feet are there in a mile? How many minutes in an hour? Dimensional analysis means take \[ 30 \frac{miles}{hour} times \frac {feet} {miles}times \frac{minutes}{hour}\] and solve. Make sure that everything cancels properly (this example includes a deliberate error)